OpenStudy (rebeccapink6398):

WILL MEDAL+FAN!!! solve for n (8 3/4+7 2/5)+6 3/5=n+14

OpenStudy (anonymous):

323/20= n +37/5 n= 35/4

OpenStudy (rebeccapink6398):

@maths911 they don't have 34/4...

OpenStudy (rebeccapink6398):

A. 9 3/5 B. 7 4/7 C. 7 3/4 D.8 3/4

OpenStudy (anonymous):

35/4 = 8 3/4 D

OpenStudy (rebeccapink6398):

Thx! I thought so :D